OpenStudy (anonymous):

Which of the following is equivalent to log(sub)40 8 rounded to three decimal places?

OpenStudy (anonymous):

On this one I would do this 40^x=8 then multiply each side by ln ln40^x=ln8 using rules of logs bring the exponent out front and divide ln40 out x=ln8/ln40 and use your cal to solve

OpenStudy (anonymous):

0.564